You require 20 gel-ink pens. STORE XXX sells these pens at R 9.50 each, while store YYY sells them at
R 10.00 each, but have a 6% sale on all items. Further, store WWW sells these pens at R 19.00 each with a buy one get one free special. Finally, store MMM sells the pens at R 14.50 each with a third off the total price.

Which store offers the best deal? Show all calculations
1
Expert's answer
2018-04-10T17:12:10-0400
Let’s calculate total prices for each shop:
XXX: R 9,50 * 20 = R 190
No discount.
Total: R 190
YYY: R 10 * 20 = R 200
6% off everything.
1% of 200 is 2 => 6% = 12
200 - 12 = 188
Total: R 188
WWW: buy one get one free means that to get 20 pens we only need to buy 10 =>
R 19 * 10 = R 190
Total: R 190
MMM: first, we buy 20 pens:
R 14.50 * 20 = R 290
And then get 1/3 off the total price:
290 / 3 = 96.7
Final price: 290 - 96.7 = 193.3
Total: R 193

The best deal for buying 20 gel-ink pens is offered by the YYY shop and equals R 188.